How much of 0. 3 AM start solution of sucrose what you need to make 150 mL dilution of sucrose at a final concentration 0.1


1
Expert's answer
2021-08-25T06:37:46-0400

Molar mass of Sucrose = 342.297

0.3 × 342.297 = 102.69

= 150 - 102.69 = 47.31ml

= 0.1 × 47.31 = 473.1ml